AOB

Speaker : All ( All )


Talk time : 12:55, Duration : 00:05
  • New Library status SL13c
  • Run 12 production status

Incomings:

  • KFParticles/HLT activities planned for July 27th - August 31st
  • Stv physics evaluation may need to be resumed soon
  • Run 12 re-production will be needed (pending understanding of the dE/dx offset issues)
  • Possible incoming farm upgrade to SL6 (new try) - TBD